Please help me solve this problem:
A class consists of 2/5 freshmen, 1/4 sophomores, and
1/10 juniors; the rest are seniors. What fraction of the class is
seniors?

Answer by ghali0079(12)

2/5 + (1/4) + (1/10)
lets write it down all /20.
8/20 + (5/4) +(2/20)
8+5+2/20 = 15/20
so senior will be 1-15/20= 5/20